Their is a tower of height 25m.At the top of the tower a man is standing with a ball in his hand.The man throws the ball with a velocity of 20m/sec in the upward direction.Find the time taken by the ball to reach the ground.

Grade:12

2 Answers

Sumit Majumdar IIT Delhi
8 years ago
From the top of the tower, the time it would take to reach the highest point would be given by:
t=\frac{u}{g}=\frac{20}{10}=2 sec
The same time would be taken to come back to the top of the tower.
Now the time it would take to come to ground from the top of the tower can be found from the formula:
s=ut+\frac{1}{2}gt^2\Rightarrow25=20\timest+\frac{1}{2}\times10\times t^2\Rightarrow t= 1 sec
Hence, the total time taken would be equal to: 2+2+1=5 sec.
Here, we havent taken the height of the man into account, as its not given.
sajid
8 years ago
s=ut+\frac{1}{2}gt^2\Rightarrow25=20\timest+\frac{1}{2}\times10\times t^2\Rightarrow t= 1 secto reach highest point t=u/g=2 sec
total time=2+2+1=5 sec